OpenStudy (anonymous):

Really stuck on this one! please help! :) Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel

OpenStudy (anonymous):

x = miles driven 75 + 2*x = 80 + 1,5*x x = 10 miles

OpenStudy (anonymous):

okay... so somehow i dont understand how i got 2.5. ill show u my work, maybe you can tell me what i did wrong...

OpenStudy (anonymous):

and thanks too. typing up my work right now

OpenStudy (anonymous):

Here is how you solve it 75 + 2*x = 80 + 1.5*x 2*x - 1.5*x = 80-75 0.5*x = 5 x = 5/0.5 x = 10

OpenStudy (anonymous):

75 + 2x = 80 + 1.5x -75 -75 2x =5+1.5x -1.5x -1.5x .5x=5 .5x/.5 = 5/.5 OH thats what i did wrong

OpenStudy (anonymous):

i did 5/2

OpenStudy (anonymous):

thank you so much!!! i guess it was a silly mistake on my part

OpenStudy (anonymous):

anytime